# Conformal Prediction for Manifold-based Source Localization with Gaussian Processes

### v1

@@ I Introduction
Sound source localization (SSL) is essential for various audio applications, including automated camera steering [1], speech enhancement and separation [2], and robot audition [3]. Numerous classical SSL methods have been developed over the last decades, including Multiple Signal Classification (MUSIC) [4], Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) [5], Time-Difference-of-Arrival (TDOA)-based approaches [6, 7], Steered Response Power (SRP) techniques [8] and clustering of time-difference-of-arrival (TDOA) estimates across time frequency bins [2, 9]. More recently, a variety of deep learning techniques have also been introduced [10, 11, 12]. However, errors are unavoidable in any localization method and can significantly affect decision-making or other systems that rely on these outcomes. For example, robots and drones depend on precise location estimates to plan their next actions, while an automated camera may need to zoom out when location estimates are uncertain. Therefore, it is essential to accurately quantify the uncertainty in localization estimates.
Conformal prediction (CP) is a versatile framework for generating prediction intervals with guaranteed coverage, ensuring that the true value falls within the PIs with a user-specified probability [13]. Recently, CP has been applied to source localization in a few studies. Previous work has explored its implementation with various uncertainty measures, such as Monte Carlo dropout, model ensembles, and quantile regression [14]. In the context of multi-source direction-of-arrival (DOA) estimation, a Gaussian mixture model was employed to parameterize the conditional multi-source DOA distribution, and CP was used to derive statistical PI from the mixture model outputs [15]. However, these studies primarily focused on synthetic anechoic data and did not incorporate real speech signals. Additionally, they employ an Inductive CP (ICP) approach, which relies heavily on a substantial amount of held-out calibration data in addition to the training data.
@@
To address these challenges, we adopt a semi-supervised approach for source localization [16, 17]. Our approach assumes access to a specific room where we can collect unlabeled data (i.e., measurements from unknown source locations) along with a small set of labeled samples with known source positions. To estimate the source positions, we employ Gaussian process regression (GPR) using a specialized manifold-based kernel that captures the geometric structure of the acoustic features, leveraging both labeled and unlabeled data. We then integrate this method with a TCP framework, which does not require additional held-out calibration data and can be efficiently applied to GPR [18]. We extensively evaluate our approach under different simulated acoustic conditions, demonstrating its effectiveness in producing valid PIs.
@@ II-A Problem Formulation
where $\hat{S}_{m,21}(k)$ and $\hat{S}_{m,11}(k)$, represents the estimated cross-spectrum between microphones and the spectrum of the first microphone, respectively, when considering the measurements of the $m$-th node. We focus on a pre-defined frequency range $k_{1},\ldots,k_{F}$, where reliable RTF estimates can be expected, and define the RTF vector $\mathbf{h}^{m}=[\bar{h}_{m}(k_{1}),\ldots,\bar{h}_{m}(k_{F})]^{T}$ associated with the $m$th node. We assume that the RTF samples of the $m$-th node reside on a low dimensional manifold, denoted as $\mathcal{M}_{m}$. This assumption is based on the fact that in a given environment, where both the acoustic conditions and microphone positions remain relatively constant, the primary variable affecting the RTF samples is the change in source position [19, 17]. Since each node has a different perspective, it will exhibit distinct relationships between the RTF samples, resulting in a unique manifold structure for each node. Finally, we define an aggregated RTF vector for all nodes by: $\mathbf{h}=[\mathbf{h}^{1},\ldots,\mathbf{h}^{m}]^{T}$.
@@ II-B Multiple Manifold Gaussian Process
where $\tilde{\mathbf{K}}_{L}$ is an $n_{L}\times n_{L}$ covariance matrix defined over the labeled samples, $\tilde{\mathbf{K}}_{L,t}$ is an $n_{L}\times 1$ covariance vector between the labelled samples and $f(\mathbf{h}_{t})$, $\tilde{\mathbf{K}}_{t}$ is the variance of $f(\mathbf{h}_{t})$, and $\mathbf{I}_{n_{L}}$ is the $n_{L}\times n_{L}$ identity matrix. Note that although the variance (10) reflects estimation uncertainty, if the model is not well-specified (e.g. wrong GP model hyperparameters or likelihood [20, 21]), these uncertainty estimates can become misleading. To mitigate this, we employ CP to produce valid PIs, even when the model is completely misspecified.
@@ II-C Conformal Prediction with Gaussian Process Regression
CP [13] is a framework that constructs prediction intervals (PIs) with a guaranteed coverage probability, defined by the user. There are two primary approaches to CP: Inductive CP (ICP)[22] and Transductive CP (TCP)[23]. ICP enhances computational efficiency by dividing the dataset into separate training and calibration subsets, while TCP leverages the entire dataset for both tasks, resulting in higher statistical efficiency. However, in TCP, the model must be re-computed for each test sample and every candidate label, which can be computationally expensive. Consequently, TCP is typically applied to specific models [24, 25], where this re-computation remains feasible. Recently, GPR-CP [18],was introduced, combining GPR with TCP, to generate PIs with guaranteed confidence levels while maintaining statistical efficiency. We will integrate this approach into the manifold-based GPR localization method.
@@
(CP coverage guarantee [13]). Suppose $(\mathbf{h}_{1},p_{1}),\ldots,(\mathbf{h}_{n_{L}},p_{n_{L}}),(\mathbf{h}_{t},p_{t})$ are exchangeable, then the PI (13) satisfies:
@@
In principal, the computation of (11) requires re-estimating the model for each held-out combination and every candidate test position. Below, we outline how the PIs can be efficiently computed with a GPR as the underlying model [18]. We utilize the fact that the LOO estimator for GPR can be written as [26]:
@@
and the interval $\Gamma_{t}^{\delta}(\mathbf{h}_{t})$ contains all candidates for which $\mathbb{P}_{\textrm{val}}(\tilde{p}_{t})>\delta$, as defined in (13). An efficient algorithm for finding all relevant intervals appears in [18, 24].
In addition, the nonconformity scores can be normalized to obtain more precise PIs, which are tighter for inputs that are easier to predict and wider for inputs that are more difficult to predict [18]. We use the LOO-variance (16) for normalization:
@@ III-A Experimental settings
We simulated $5.2\times 6.2\times 3.5$m rooms with reverberation times of $300$ ms and $700$ ms. Five ($M=5$) microphone pairs were placed around the room’s walls. The source positions were confined to a $2\times 2$m squared region (see Fig. 1). We generated $n_{L}=225$ labeled samples, equally spaced with $0.133$m resolution, $n_{U}=100$ unlabeled samples and $n_{T}=200$ test samples, both randomly sampled within the same region. Simulated room impulse responses [27] were convolved with white Gaussian noise signals to produce the labeled measurements, while speech utterances from the LibriSpeech dev-set [28] were used for the unlabeled and test samples. White Gaussian noise was added to all microphone signals, with an SNR of $15$dB. The signals were analyzed using an STFT with $1024$ frequency bins and $75\%$ frame overlap. The frequency range $150-1500$Hz was used in construction of the RTF vectors. GPyTorch [29] was employed for optimizing the model’s parameters $\Theta=\{\sigma_{1},\ldots,\sigma_{M},\sigma_{p}\}$. We set $\gamma=8$ for the score normalization (22).

### v2

@@ I Introduction
Sound source localization (SSL) is essential for various audio applications, including automated camera steering [huang2000passive], speech enhancement and separation [mandel2010model], and robot audition [rascon2017localization]. Numerous classical SSL methods have been developed over the last decades, including Multiple Signal Classification (MUSIC) [schmidt1986multiple], Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) [roy1989esprit], Time-Difference-of-Arrival (TDOA)-based approaches [brandstein1997closed, schau1987passive], Steered Response Power (SRP) techniques [dibiase2000high] and clustering of time-difference-of-arrival (TDOA) estimates across time frequency bins [mandel2010model, schwartz2013speaker]. More recently, a variety of deep learning techniques have also been introduced [opochinsky2021deep, grinstein2023neural, grumiaux2022survey]. However, errors are unavoidable in any localization method and can significantly affect decision-making or other systems that rely on these outcomes. For example, robots and drones depend on precise location estimates to plan their next actions, while an automated camera may need to zoom out when location estimates are uncertain. Therefore, it is essential to accurately quantify the uncertainty in localization estimates.
Conformal prediction (CP) is a versatile framework for generating prediction intervals with guaranteed coverage, ensuring that the true value falls within the PIs with a user-specified probability [vovk2005algorithmic]. Recently, CP has been applied to source localization in a few studies. Previous work has explored its implementation with various uncertainty measures, such as Monte Carlo dropout, model ensembles, and quantile regression [khurjekar2023uncertainty]. In the context of multi-source direction-of-arrival (DOA) estimation, a Gaussian mixture model was employed to parameterize the conditional multi-source DOA distribution, and CP was used to derive statistical PI from the mixture model outputs [khurjekar2024multi]. However, these studies primarily focused on synthetic anechoic data and did not incorporate real speech signals. Additionally, they employ an Inductive CP (ICP) approach, which relies heavily on a substantial amount of held-out calibration data in addition to the training data.
@@
To address these challenges, we adopt a semi-supervised approach for source localization [laufer2016multiple, laufer2020data]. Our approach assumes access to a specific room where we can collect unlabeled data (i.e., measurements from unknown source locations) along with a small set of labeled samples with known source positions. To estimate the source positions, we employ Gaussian process regression (GPR) using a specialized manifold-based kernel that captures the geometric structure of the acoustic features, leveraging both labeled and unlabeled data. We then integrate this method with a TCP framework, which does not require additional held-out calibration data and can be efficiently applied to GPR [papadopoulos2024guaranteed]. We extensively evaluate our approach under different simulated acoustic conditions, demonstrating its effectiveness in producing valid PIs.
@@ II-A Problem Formulation
where $\hat{S}_{m,21}(k)$ and $\hat{S}_{m,11}(k)$, represents the estimated cross-spectrum between microphones and the spectrum of the first microphone, respectively, when considering the measurements of the $m$-th node. We focus on a pre-defined frequency range $k_{1},\ldots,k_{F}$, where reliable RTF estimates can be expected, and define the RTF vector $\mathbf{h}^{m}=[\bar{h}_{m}(k_{1}),\ldots,\bar{h}_{m}(k_{F})]^{T}$ associated with the $m$th node. We assume that the RTF samples of the $m$-th node reside on a low dimensional manifold, denoted as $\mathcal{M}_{m}$. This assumption is based on the fact that in a given environment, where both the acoustic conditions and microphone positions remain relatively constant, the primary variable affecting the RTF samples is the change in source position [laufer2015study, laufer2020data]. Since each node has a different perspective, it will exhibit distinct relationships between the RTF samples, resulting in a unique manifold structure for each node. Finally, we define an aggregated RTF vector for all nodes by: $\mathbf{h}=[\mathbf{h}^{1},\ldots,\mathbf{h}^{m}]^{T}$.
@@ II-B Multiple Manifold Gaussian Process
where $\tilde{\mathbf{K}}_{L}$ is an $n_{L}\times n_{L}$ covariance matrix defined over the labeled samples, $\tilde{\mathbf{K}}_{L,t}$ is an $n_{L}\times 1$ covariance vector between the labelled samples and $f(\mathbf{h}_{t})$, $\tilde{\mathbf{K}}_{t}$ is the variance of $f(\mathbf{h}_{t})$, and $\mathbf{I}_{n_{L}}$ is the $n_{L}\times n_{L}$ identity matrix. Note that although the variance (10) reflects estimation uncertainty, if the model is not well-specified (e.g. wrong GP model hyperparameters or likelihood [assael2014heteroscedastic, griffiths2021achieving]), these uncertainty estimates can become misleading. To mitigate this, we employ CP to produce valid PIs, even when the model is completely misspecified.
@@ II-C Conformal Prediction with Gaussian Process Regression
CP [vovk2005algorithmic] is a framework that constructs prediction intervals (PIs) with a guaranteed coverage probability, defined by the user. There are two primary approaches to CP: Inductive CP (ICP)[papadopoulos2002inductive] and Transductive CP (TCP)[Vovk2015TransductiveCP]. ICP enhances computational efficiency by dividing the dataset into separate training and calibration subsets, while TCP leverages the entire dataset for both tasks, resulting in higher statistical efficiency. However, in TCP, the model must be re-computed for each test sample and every candidate label, which can be computationally expensive. Consequently, TCP is typically applied to specific models [nouretdinov2001ridge, papadopoulos2011regression], where this re-computation remains feasible. Recently, GPR-CP [papadopoulos2024guaranteed],was introduced, combining GPR with TCP, to generate PIs with guaranteed confidence levels while maintaining statistical efficiency. We will integrate this approach into the manifold-based GPR localization method.
@@
(CP coverage guarantee [vovk2005algorithmic]). Suppose $(\mathbf{h}_{1},p_{1}),\ldots,(\mathbf{h}_{n_{L}},p_{n_{L}}),(\mathbf{h}_{t},p_{t})$ are exchangeable, then the PI (13) satisfies:
@@
In principal, the computation of (11) requires re-estimating the model for each held-out combination and every candidate test position. Below, we outline how the PIs can be efficiently computed with a GPR as the underlying model [papadopoulos2024guaranteed]. We utilize the fact that the LOO estimator for GPR can be written as [rasmussen2006gaussian]:
@@
and the interval $\Gamma_{t}^{\delta}(\mathbf{h}_{t})$ contains all candidates for which $\mathbb{P}_{\textrm{val}}(\tilde{p}_{t})>\delta$, as defined in (13). An efficient algorithm for finding all relevant intervals appears in [papadopoulos2024guaranteed, nouretdinov2001ridge].
In addition, the nonconformity scores can be normalized to obtain more precise PIs, which are tighter for inputs that are easier to predict and wider for inputs that are more difficult to predict [papadopoulos2024guaranteed]. We use the LOO-variance (16) for normalization:
@@ III-A Experimental settings
We simulated $5.2\times 6.2\times 3.5$m rooms with reverberation times of $300$ ms and $700$ ms. Five ($M=5$) microphone pairs were placed around the room’s walls. The source positions were confined to a $2\times 2$m squared region (see Fig. 1). We generated $n_{L}=225$ labeled samples, equally spaced with $0.133$m resolution, $n_{U}=100$ unlabeled samples and $n_{T}=200$ test samples, both randomly sampled within the same region. Simulated room impulse responses [habets2006rirgenerator] were convolved with white Gaussian noise signals to produce the labeled measurements, while speech utterances from the LibriSpeech dev-set [panayotov2015librispeech] were used for the unlabeled and test samples. White Gaussian noise was added to all microphone signals, with an SNR of $15$dB. The signals were analyzed using an STFT with $1024$ frequency bins and $75\%$ frame overlap. The frequency range $150-1500$Hz was used in construction of the RTF vectors. GPyTorch [gardner2018gpytorch] was employed for optimizing the model’s parameters $\Theta=\{\sigma_{1},\ldots,\sigma_{M},\sigma_{p}\}$. We set $\gamma=8$ for the score normalization (22).